062818 UCSC fog update

Although earlier this week there was more fog, today no viable sample was collected. The collector was cleaned yesterday. I took a blank sample of 160ml and then treated it with HCl.

Took more acid cleaned jars to roof

-Marcella

Leave a Reply

Your email address will not be published. Required fields are marked *